A line contains the point ( - 6, 7) and has a slope of . Write an equation of the line satisfying the given conditions. Write the answer in slope-intercept form.

and

A line contains the point ( - 3, - 4) and is perpendicular to a line with a slope of . Write an equation of the line satisfying the given conditions. Write the answer in slope-intercept form.

Found 3 solutions by ikleyn, greenestamps, josgarithmetic:
Answer by ikleyn(52781)

.

            Read your post and try to understand why it is nonsense.


Hey, one more such a post, and I will write a MAIL to your managers asking them to fire you from your position.

--------------
Comment from student: It can't be nonsense if its an actual question from school. its multiple choice but i refer to see how its broken down. These are the answers to choose from. If you can't help please do stop others from helping. A line contains the point ( - 3, - 4) and is perpendicular to a line with a slope of . Write an equation of the line satisfying the given conditions. Write the answer in slope-intercept form. A. y=-5/3x-4 B. y=-5/3x-9 C. y= 5/3x-9 D. y=-5/3x-3
--------------



MY response :


        Read your post and try to understand why it is nonsense.


                        Read your post and try to understand why it is nonsense.


                                        Read your post and try to understand why it is nonsense.




Answer by greenestamps(13200)



To the student....

Since you continue to post the questions in the same way, and one of the tutors here keeps telling you it is nonsense the way you post them, I can only guess that you don't know how the questions appear on the forum.

In EVERY case, the problem talks about "... a slope of ." No number is given for the slope in every case. Are you using some strange font or something like that which causes the slopes to not show up in your post? I would hope that is the case, since you surely know that the questions can't be answered without knowing the slope of the given line.

Answer by josgarithmetic(39617)

Another question like this was posted (MISSING THE SLOPE) yesterday, and it was answered. Look at THAT one and understand how it was solved.